OpenStudy (anonymous):

HELP IMA GONNA FAIL IF I DONT GET THESE RIGHT SO HELP ME WILL FAN AND MEDAL!!!!!!

OpenStudy (anonymous):

The graph below shows the relationship between pounds of dog food and total cost, in dollars, for the dog food. A coordinate graph is shown. The label on the x axis is Weight in pounds. The label on the y axis is Total cost in dollars. There is a line connecting the points 0, 0 and 10, 5. Select each statement about the graph that is true. A. Point (0, 0) shows the cost is $0 for 0 pound of dog food. B. Point (1, 0.5) shows the cost is $1.00 for 0.5 pound of dog food. C. Point (2, 1) shows that 2 pounds of dog food cost $1. D. Point (4, 2) shows the cost is $4 for 2 pounds of dog food. E. Point (10, 5) shows that 10 pounds of dog food costs $5.

OpenStudy (anonymous):

@iGreen

jabez177 (jabez177):

@Arianna842 @AriPotta can you help her please

OpenStudy (anonymous):

I want to say its A but not sure @ahvei Can you help her

OpenStudy (anonymous):

its multiple

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

can you quit spamming? thanks

OpenStudy (anonymous):

then I have no clue @ahvei @AMYCARTER @Abhisar @gummybear1021 @Hannah15121 then help her MONSTER

OpenStudy (anonymous):

Im not sure, but if i had to guess i would say A.... dont hate me if im wrong xD

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

or E.. idek lol

OpenStudy (aripotta):

ok so you're given two points, (0,0) and (10,5). since i can't see the graph, we're going have to make the equation of the line. using the points and the slope formula, the slope is 5/10 or 1/2. so the equation is y = 1/2x

OpenStudy (aripotta):

we're going to use that equation for each choice to see if they're true. we know for sure that A is correct, since (0,0) is one of the given points

OpenStudy (ultrasoundwave15):

I think I'd go with A

OpenStudy (aripotta):

b is (1,0.5) so 0.5 = 1(1/2) which is true but it says "$1.00 for 0.5 pound" which is false since the x axis is weight c is (2,1) so 1 = 2(1/2) which is true, which is 2 pounds for $1. so c is correct d is (4,2) so 2 = 4(1/2) which is true, which is 4 pounds for $2, so d is not correct e is (10,5) so 5 = 10(1/2) which is true, which is 10 pounds for $5, so e is also correct

OpenStudy (aripotta):

so a, c, and e are your answers

OpenStudy (anonymous):

thx let me show you the next question please

OpenStudy (aripotta):

can you make a new question?

OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $ Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in by

OpenStudy (anonymous):

im not sure on this one :/

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (aripotta):

ok so let's say september is s. 20% increase can be put as 120% total. 120% = 1.2, so 1.5 = 1.2s solve for s

OpenStudy (anonymous):

I already know that part 1.20

OpenStudy (anonymous):

but how do I solve for S?

OpenStudy (aripotta):

since s is being multiplied by 1.2, divide both sides by 1.2 because division is the opposite of multiplication

OpenStudy (anonymous):

1.25

OpenStudy (aripotta):

yep. so that was the price in september

OpenStudy (anonymous):

so which moth did they make more money in and by how much

OpenStudy (ultrasoundwave15):

I believe it is E as well

OpenStudy (aripotta):

The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. ok so for september, you would do 0.4(900 x 1.25) then for october, you would do 0.4(700 x 1.5)

OpenStudy (anonymous):

ok so skip that cus I already know that

OpenStudy (aripotta):

the multiplication in the parentheses is going to give you how much money they made for that many flowers. multiplying the 0.4 will give you how much money they made, when they only earn 40% for each flower, since 40% = 0.4

OpenStudy (anonymous):

oh so its september by 30

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

A car travels 1 6 of the distance between two cities in 3 5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour? Enter your fraction in the space provided.

OpenStudy (aripotta):

can you make a new question please

OpenStudy (aripotta):

and maybe give me a medal? :3

OpenStudy (anonymous):

I will give you a medal when you answer all the questions

OpenStudy (anonymous):

so answer them please

OpenStudy (aripotta):

you're not supposed to put multiple questions in a thread. and how is it fair that i would only get 1 medal for helping you with 5 or more questions

OpenStudy (anonymous):

well this is three so your good lol

OpenStudy (anonymous):

tell you what Ill will make the new question for the last question ok?

OpenStudy (aripotta):

how many are there

OpenStudy (anonymous):

4 were on the third

OpenStudy (aripotta):

ok, i'll do it for now. but for future reference, please post each question in their own thread :)

OpenStudy (anonymous):

ok well do this question for me A car travels 1 6 of the distance between two cities in 3 5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour? Enter your fraction in the space provided.

OpenStudy (aripotta):

make a proportion\[\frac{ distance }{ time }\frac{ \frac{ 1 }{ 6 } }{ \frac{ 3 }{ 5 } }=~\frac{ x }{ 1 }\] cross-multiply and solve for x

OpenStudy (anonymous):

how do i do that again

OpenStudy (aripotta):

1/6(1) = 3/5x 1/6 = 3/5x

OpenStudy (anonymous):

wait no I got it is it 0.10

OpenStudy (anonymous):

?

OpenStudy (aripotta):

\[\frac{ 1 }{ 6 }\div \frac{ 3 }{ 5 }=~\frac{ 1 }{ 6 }\times \frac{ 5 }{ 3 }=~\frac{ 5 }{ 18 }\]

OpenStudy (anonymous):

if you want your medal you gotta do 13 more lol you just got trollled

OpenStudy (aripotta):

lol. bye.

OpenStudy (ultrasoundwave15):

@jabez177 remember me lol. It's Danny... this is my old old account xDD

jabez177 (jabez177):

Ikrrrr!!! I seriously never knew you were UltraSoundwave15!!! :D